OpenStudy (anonymous):

Find a ratio equivalent to 72/81 thenuse the ratio to find a proportion.

OpenStudy (anonymous):

\[\frac{ 72 }{ 81 }=\frac{ 8 }{ 9 } ,...\left( divide~the~numerator~and~denominator ~by~9 \right)\] 8 : 9